OpenStudy (anonymous):

77(x-6) = 42(2x-14) x= _____

OpenStudy (anonymous):

77(x-6) = 42(2x-14) 77x-462 = 84x-588 Subtract 77x from both sides -462 = 7x-588 Add 588 to both sides 126 = 7x Divide 7 from both sides 18 = x

OpenStudy (anonymous):

Thanks.

OpenStudy (anonymous):

I would start by dividing both sides by 7 to keep the numbers smaller....